Reports 1-1 of 1 Clear search Modify search
Detector Characterisation (Glitches)
mwas - 14:32 Wednesday 27 August 2025 (67574) Print this report
New type of glitches around 33Hz

Figure 1. A new family of glitches started on Sunday around 15:30 UTC. These glitches are loud and frequent, and look to be around 33Hz.

Figure 2. These glitches are continuing during last night, so this problem is still present. 

Images attached to this report
Comments to this report:
narnaud - 14:46 Wednesday 27 August 2025 (67576) Print this report
fiori - 14:59 Wednesday 27 August 2025 (67577) Print this report

These glitches (I guess these are the same... and 33 Hz or so might be the energy baricenter?) seems to have a precise periodic pattern, see the attached.

Images attached to this comment
salvador - 15:03 Wednesday 27 August 2025 (67578) Print this report

Their frequency distribution seems to go from around 15 Hz up to almost 60 Hz.

First glitch is most likely a 25-min glitch. (Picture shows the whitened spectrogram of Hrec_hoft_20000Hz)

Images attached to this comment
direnzo - 15:18 Wednesday 27 August 2025 (67580) Print this report

These glitches look similar to something we’ve already seen in April 2024: #64114.

fiori, paoletti, dattilo - 15:44 Wednesday 27 August 2025 (67581) Print this report

The time pattern and magnetic witness point to a noise from the power supply of the heating belts, as evidenced by Bas in O3:  https://logbook.virgo-gw.eu/virgo/?r=48521

sorry I missed the previous post, pointing to the same problem. The time pattern is clear.

However, consulting with Federico and Vincenzo: the cause is likely the power supply, whose fans are old and do not cool down enough, the power supply goes in protection ad switches on/off.

The suggestion is  first inspect and then replace.

Images attached to this comment
dattilo, zaza - 17:56 Wednesday 27 August 2025 (67582) Print this report
As a follow up of the recent investigations on this new type of glitches, we made a check of the WI heating belt power supply.
At around 14:57 UTC, with ITF mode set to Adjusting, we went in TCS room where the power supply is located, and also looking at its analog display, we saw that it was delivering intermittently, as indicated by the LSC_WI_HB_moni signal.
We replaced the power supply with a spare one of the same kind (Kert 420). During the intervention the operator temporarily opened the WI etalon loop. It was turned back on at 15:25 UTC.
Now the loop seems working regularly (attached fig).
Unfortunately, during the intervention the ITF unlocked, just a few seconds after the etalon loop was turned back on.
Images attached to this comment
salvador - 9:39 Thursday 28 August 2025 (67585) Print this report
dattilo - 12:51 Thursday 28 August 2025 (67586) Print this report
A detail about the power supply that was malfunctioning: the problem was related to its cooling fan, which was no longer working. This was causing the driver to overheat, so after about 130 seconds the protection thermostat would intervene, blocking the power delivery. At that point, the driver would cool down, and after about 5 minutes, it would start supplying power again. Then the cycle would start again.
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×